OpenStudy (anonymous):

"3 is what percent of 65?" Which of the following is the most reasonable answer to this question?

OpenStudy (anonymous):

To solve any percentage problem, you divide the number you're trying to find the percent of, by the number you're given. In this case, you divide 3 by 65. So, 3/65 = .05(rounded) which = 5% You can test this out with more simple fractions. Like, say you want to find, "1 is what percent of 4", do 1/4 which is equal to .25 which = 25%

OpenStudy (anonymous):

"is" means equals, "what" means x "of" means multiplication. 3 is what percent of 65 means 3 = x % 65

OpenStudy (anonymous):

Let the percentage be x: (x/100)*65 = 3 65x/100 = 3 x = (3*100)/65 x = 300/65 Divide it and get the answer...

OpenStudy (anonymous):

21

OpenStudy (anonymous):

It's certainly not 21. If that's the answer you got.